# bibliography.bib

### v3

@@ Abstract
This paper presents how to use common random number (CRN) simulation to evaluate Markov chain Monte Carlo (MCMC) convergence to stationarity. We provide an upper bound on the Wasserstein distance of a Markov chain to its stationary distribution after $N$ steps in terms of averages over CRN simulations. We apply our bound to Gibbs samplers on a variance component model, a model related to James-Stein estimators, and a Bayesian linear regression model. For the former two examples, we show that the CRN simulated bound converges to zero significantly more quickly compared to available drift and minorization bounds.
@@ 1 Introduction
Markov chain Monte Carlo (MCMC) algorithms are often used to simulate from a stationary distribution of interest (see e.g. [intromcmc]). One of the primary questions when using these Markov chains is, after how many iterations is the distribution of the Markov chain sufficiently close to the stationary distribution of interest, i.e. when should actual sampling begin [honestexp]. The number of iterations it takes for the distribution of the Markov chain to be sufficiently close to stationarity is called the burn-in period. Various informal methods are available for estimating the burn-in period, such as effective sample size estimation, the Gelman-Rubin diagnostic, and visual checks using traceplots or autocorrelation graphs [glm, mode, bayemethfin, convdiag]. However, none of these methods provide a formal estimate of the distance between the distribution of the Markov chain and the stationary distribution.
@@
From a theoretical perspective, distance to stationarity is traditionally measured in terms of total variation distance (e.g. [probsurv, tierney1994markov]), though more recently the Wasserstein distance has been considered [gibbswass, waslinearmixedmodels, wasandoneshot, wasmeth]. However, finding upper bounds on either distance can be quite difficult to establish [intromcmc, honestexp], and if an upper bound is known, it is usually based on complicated problem-specific calculations [waslinearmixedmodels, geomconvrates, sixta, localcont]. This motivates the desire to instead estimate convergence bounds from actual simulations of the Markov chain, which we consider here.
@@
One common method for generating upper bounds on the Wasserstein distance is through a contraction condition ([sixta, avcont]). This can often be established using common random number (CRN) simulation, i.e. using the same sequence of random variables to simulate two copies of a Markov chain with different initial values. This simulation method is also sometimes referred to as synchronous coupling [sdesynch, browniansynch]. Estimating Markov chain convergence rates using CRN simulation was first proposed in [crnconv] to find estimates of mixing times in total variation distance. This approach falls under the general framework of “auxiliary simulation” [simconv], i.e. using extra preliminary Markov chain runs to estimate the convergence time needed in the final run. More recently, [biswas] showed how CRN simulation could be used for estimating an upper bound on the Wasserstein distance (their Proposition 3.1), and provided useful applications of the CRN method to high-dimensional and tall data (their Section 4). CRN simulation to estimate convergence rates has been used on a wide range of applications including Langevin algorithms [langevin], stochastic gradient descent [sgd], stochastic differential equations (SDE) [sdesynch], and Hamiltonian Monte Carlo [hmc, unbiasedhmc].
@@
The optimality of CRN simulation is promising. In this paper we show that for our real data examples, common random number bounds perform significantly better than drift and minorization (DnM) bounds (see figures 4 and 6 for graphical comparisons). Under certain conditions, [browniansynch] shows that two reflected Brownian motion processes simulated using CRN converge to 0 almost surely, [sdesynch] shows that CRN simulation is the optimal solution for the Wasserstein distance between the solutions of SDEs, [crn] shows that CRN simulation is the joint distribution that minimizes the expected Euclidean distance, and [independentmetropolishastings] shows that CRN simulation on the independent Metropolis algorithm is an unbiased estimate of the total variation distance.
@@
In Theorem 3.3, using CRN simulation, we provide an estimated upper bound between the Wasserstein distance of a Markov chain and the corresponding stationary distribution when only the unnormalized density of the stationary distribution is known.
@@
This paper is organized as follows. In Section 2, we present definitions and notation. In Section 3, we establish convergence bounds of a Markov chain to its corresponding stationary distribution using CRN simulation when the initial distribution is not in stationarity. In Section 4 we apply our Theorem 3.3 to a Gibbs sampler for a variance component model and compare the DnM bound in Real Data Example 4.2. In Section 5, we apply our Theorem 3.3 to a Gibbs sampler for a model related to the James-Stein estimator and compare the CRN bound to a DnM and one-shot coupling bound in Real Data Example 5.1. In Section 6, we apply our Theorem 3.3 to a Gibbs sampler for a Bayesian linear regression model.
@@ 2.1 Notation
Let $(\mathcal{X},\mathcal{F})$ be a measurable space with metric $d$. The $L^{p}-$Wasserstein distance between $\pi,\nu$ on $(\mathcal{X},\mathcal{F})$ is defined as, $W_{d,p}(\pi,\nu)=\inf_{X\sim\pi,Y\sim\nu}E[d(X,Y)^{p}]^{1/p}$. Total variation is defined as $||\pi-\nu||_{TV}=\inf_{X\sim\pi,Y\sim\nu}P(X\neq Y)=\sup_{A\in\mathcal{F}}|\pi(A)-\nu(A)|$. If $\mathcal{X}\subset\mathbb{R}^{q}$, $q\in\mathbb{N}$, define the norm $||x||_{p}=(\sum_{i=1}^{q}|x_{i}|^{p})^{1/p}$ where $p\in[1,\infty)$. For a distribution $\nu$, the density function is denoted as $f_{\nu}$. If $X\perp\!\!\!\perp Y$, then $X$ and $Y$ are independent.
@@
Define a Markov chain $\{X_{n}\}_{n\geq 1}$ in $\mathcal{X}$ such that $X_{n}=k_{\theta_{n}}(X_{n-1})$, where $\{\theta_{n}\}_{n\geq 1}$ are i.i.d. random variables on some measurable space $\Theta$ and random measurable mappings $k_{\theta_{n}}:\mathcal{X}\mapsto\mathcal{X}$. The set of random functions $k_{\theta_{1}},k_{\theta_{2}},\ldots$ is called an iterated function system. Any time-homogeneous Markov chain can be represented as an iterated function system [avcont]. CRN simulation in this context occurs when we simulate two Markov chains $\{X_{n}\}_{n\geq 1}$ and $\{Y_{n}\}_{n\geq 1}$ with different initial values $x_{0}\neq y_{0}$, but with the same transition functions. That is, $\theta_{n}$ are the same in $X_{n}=k_{\theta_{n}}(X_{n-1})$ and $Y_{n}=k_{\theta_{n}}(Y_{n-1})$, $n\geq 1$. See Algorithm 1 for more details. The stationary distribution of $X_{n}$ is defined as $\mathcal{L}(X_{\infty})=\pi$.
@@
$IG^{-1}(\alpha,\beta)$ represents the inverse gamma distribution with shape parameter $\alpha$ and rate parameter $\beta$. The density function for the inverse gamma is denoted as $f_{IG}(x\mid\alpha,\beta)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}x^{-\alpha-1}e^{-\beta/x}$. $N(\mu,\sigma^{2})$ represents the normal distribution with mean $\mu$ and variance $\sigma^{2}$. The density function for the normal distribution is denoted as $f_{N}(x\mid\mu,\sigma^{2})=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}$.
Essential supremum and infimum
Let $f:\mathcal{X}\to\mathbb{R}$ be a function and $\lambda$ be the Lebesgue measure. The essential supremum is the smallest value $a\in\mathbb{R}$ such that $\lambda(x\mid f(x)<a)=1$. More formally, $ess\sup_{x}f(x)=\inf_{a\in\mathbb{R}}\{a\mid\lambda(x\mid f(x)>a)=0\}$. The essential infimum is likewise the largest value $a\in\mathbb{R}$ such that $\lambda(x\mid f(x)>a)=1$. Or, $ess\inf_{x}f(x)=\sup_{a\in\mathbb{R}}\{a\mid\lambda(x\mid f(x)<a)=0\}$ [esssup].
2.2 Rejection rate and separation distance
Definition 2.1 (Rejection sampling).
Suppose that we have a target distribution $\pi$, which we want to sample from, but is difficult to do, and we have a proposal distribution $\nu$ that is easier to sample from. Suppose also that for the corresponding density functions $f_{\pi}(x)$ and $f_{\nu}(x)$, $f_{\nu}(x)=0\implies f_{\pi}(x)=0$, where $x\in\mathcal{X}$ and that $K^{\prime}=ess\sup_{x\in\mathcal{X}}f_{\pi}(x)/f_{\nu}(x)$. To generate a random variable $X_{\infty},\mathcal{L}(X_{\infty})=\pi$ we do the following,
  1. 1.
    Sample $X\sim\nu$ and $U\sim Unif(0,1)$ independently.
  2. 2.
    If $U\leq\frac{1}{K^{\prime}}\frac{f_{\pi}(X)}{f_{\nu}(X)}$ then accept $X$ as a draw from $\pi$. Otherwise reject $X$ and restart from step 1.
Lemma 2.1 (Rejection sampler rejection rate).
For the rejection sampler algorithm defined above, denote the event $A=\{X\text{ is accepted as a draw from }\pi\}$. The rejection rate is $r(\pi,\nu)=1-P(A)=1-1/K^{\prime}$ where $K^{\prime}=ess\sup_{x\in\mathcal{X}}f_{\pi}(x)/f_{\nu}(x)$.
Proof.
See Section 11.2.2 of [ross]. ∎
We further define the separation distance on the continuous state space $\mathcal{X}$ as follows. Separation distance was first defined in [ALDOUS198769] for discrete state spaces. As far as we know, separation distance was only recently defined on a continuous state space in [sepdist]. We use the definition of separation distance defined in [sepdist] where the density functions are known.
Definition 2.2 (Separation distance (Remark 5 of [sepdist])).
Lemma 2.2.
$s(\pi,\nu)=r(\pi,\nu)$
Proof.
$=1-\frac{1}{ess\sup_{x\in\mathcal{X}}\frac{f_{\pi}(x)}{f_{\nu}(x)}}=1-\frac{1}{K^{\prime}}=r(\pi,\nu)$
3 Estimating Markov chain convergence rates using Common Random Number Simulation
We propose bounding the $L^{p}-$Wasserstein distance between the $n$th iteration of a Markov chain $\{X_{n}\}_{n\geq 1}$ and the corresponding stationary distribution $\pi$ through CRN simulation. Our main result is Theorem 3.3. Using CRN simulation as a convergence diagnostic tool was first discussed in [crnconv] for bounding total variation distance.
Before providing a method for simulating an upper bound on $W_{d,p}(\mathcal{L}(X_{n}),\pi)$, we must first provide a method of bounding the expected distance between a Markov chain $X_{n}$ and the corresponding stationary random variable $X_{\infty}$, $E[d(X_{n},X_{\infty})^{p}]^{1/p}$, given the expected distance between two random variables, $E[d(X_{n},Y_{n})^{p}]^{1/p}$.
To generate this bound we use the rejection sampler. The use of rejection sampling to generate an upper bound on the expected distance between a proposal density function $\nu$ and a stationary density function $\pi$ was motivated by [crnconv], which used rejection sampling to generate similar upper bounds in total variation distance.
@@ 3 Estimating Markov chain convergence rates using Common Random Number Simulation
Let $\{X_{n}\}_{n\geq 1}$ and $\{Y_{n}\}_{n\geq 1}$ be two copies of a Markov chain with initial distribution $\mathcal{L}(Y_{0})=\nu$, $X_{0}\perp\!\!\!\perp Y_{0}$. Assume that the corresponding stationary distribution, $\pi$, and $\nu$ are defined on the same support. Then for $p\in[1,\infty)$
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq\left(KE[d(X_{n},Y_{n})^{p}]\right)^{1/p}$ (1)
where $K\geq K^{\prime}=\frac{1}{1-s(\pi,\nu)}=\frac{1}{1-r(\pi,\nu)}$.
Proof.
Let $X_{n},Y_{n}$ be two copies of the Markov chain and $A=\{\text{Accept $Y_{0}$ as a draw from }\pi\}$. Recall $r(\pi,\nu)=1-P(A)$. Note that $\mathcal{L}(Y_{0}|A)=\pi$ and so, $\mathcal{L}(Y_{n}|A)=\pi$.
$\geq E[d(X_{n},X_{\infty})^{p}]P(A)$
$\geq E[d(X_{n},X_{\infty})^{p}]\frac{1}{K^{\prime}}$
Further, $1/K^{\prime}=1-r(\pi,\nu)=1-s(\pi,\nu)$ by Lemma 2.2. Finally, $K\geq K^{\prime}$ and $E[d(X_{n},X_{\infty})^{p}]\geq W_{d,p}(\mathcal{L}(X_{n}),\pi)^{p}$, so equation 1 follows. ∎
In many cases, the value of $K^{\prime}$ as defined in Lemma 2.1 and Theorem 3.1 is unknown. We bound the value of $K^{\prime}$ as follows.
Lemma 3.2.
Suppose that the normalizing constant for $f_{\pi}$ is unknown and that we only know of a function $g(x)$ such that $cg(x)=f_{\pi}(x)$ where $c$ is unknown. If we can generate a $K\in\mathbb{R}_{+}$ such that for $B\subset\mathcal{X}$
[…]
Then, $W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq\left(KE[d(X_{n},Y_{n})^{p}]\right)^{1/p}$.
Proof.
[…]
Remark.
Applying $d(X,Y)=I_{X\neq Y}$ to Theorem 3.3 bounds total variation distance between a Markov chain and the corresponding stationary distribution as follows,
$||\mathcal{L}(X_{n})-\pi||_{TV}=\inf E[I_{X_{n}\neq X_{\infty}}]=W_{I,1}(\mathcal{L}(X_{n}),\pi)\leq KE[I_{X_{n}\neq Y_{n}}]=KP(X_{n}\neq Y_{n})$ (3)
For stopping time $\tau=\min\{n:d(X_{n},Y_{n})=0\}$, $||\mathcal{L}(X_{n})-\pi||_{TV}\leq KP(\tau>n)$
Next we define Algorithm 1 to generate an estimate of $E[d(X_{n},Y_{n})^{p}]$ using common random number. This algorithm is similar to Algorithm 1 in [biswas].
Algorithm 1 An estimate of $E[d(X_{N},Y_{N})^{p}]\approx\frac{1}{M}\sum_{i=1}^{M}d(x_{N,i},y_{N,i})^{p}$ using CRN
@@
Let $\{X_{n}\}_{n\geq 1}$ and $\{Y_{n}\}_{n\geq 1}$ be two copies of a Markov chain with stationary distribution $\pi$ simulated according to Algorithm 1 and with initial value $\mathcal{L}(Y_{0})=\nu$ and $Y_{0}\perp\!\!\!\perp X_{0}$. Suppose that for fixed $p\in(1,\infty]$ and each $n\geq 0$, $E[d(X_{n},Y_{n})^{p}]<\infty$. Suppose also that $\nu$ and $\pi$ are defined on the same support. Then
$W_{d,p}(\mathcal{L}(X_{N}),\pi)\leq\left(K\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}d(X_{N,i},Y_{N,i})^{p}\right)^{1/p}$ (4)
holds almost surely where $K\geq K^{\prime}=\frac{1}{1-s(\pi,\nu)}=\frac{1}{1-r(\pi,\nu)}$.

### v4

# Estimating MCMC convergence rates using common random number simulations

@@ Abstract
This paper presents how to use common random number (CRN) simulation to evaluate Markov chain Monte Carlo (MCMC) convergence to stationarity. We provide an upper bound on the Wasserstein distance of a Markov chain to its stationary distribution after $N$ steps in terms of averages over CRN simulations. We apply our bound to Gibbs samplers on a model related to James-Stein estimators, a variance component model, and a Bayesian linear regression model. For the first two examples, we show that the CRN simulated bound converges to zero significantly more quickly compared to available drift and minorization bounds.
@@ 1 Introduction
Markov chain Monte Carlo (MCMC) algorithms are often used to simulate from a stationary distribution of interest (see e.g. [19]). One of the primary questions when using these Markov chains is, after how many iterations is the distribution of the Markov chain sufficiently close to the stationary distribution of interest, i.e. when should actual sampling begin [24]. The number of iterations it takes for the distribution of the Markov chain to be sufficiently close to stationarity is called the burn-in period. Various informal methods are available for estimating the burn-in period, such as effective sample size estimation, the Gelman-Rubin diagnostic, and visual checks using traceplots or autocorrelation graphs [16, 26, 41, 50]. While these methods “can detect problems with an MCMC simulation, […] they cannot prove that the simulation is generating a representative sample.” [37]
@@
Convergence analysis studies the distance of a Markov chain to stationarity and is traditionally measured in terms of total variation distance (e.g. [43, 55]), though more recently the Wasserstein distance has been considered [20, 29, 35, 40]. However, finding upper bounds on either distance can be quite difficult to establish [19, 24, 37], and if an upper bound is known, it is usually based on complicated problem-specific calculations [29, 38, 51, 52]. This motivates the desire to instead estimate convergence bounds from actual simulations of the Markov chain, which we consider here via common random number (CRN) simulation.
@@
CRN simulation occurs when we initiate two copies of the chain with different values, but use the same sequence of random variables to simulate a specific random function representation of the chain (See Section 2 for a formal definition). This simulation method is also sometimes referred to as synchronous coupling [3, 9, 28] and falls under the general framework of “auxiliary simulation” [11], i.e. using extra preliminary Markov chain runs to estimate the convergence time needed in the final run. Estimating Markov chain convergence rates using CRN simulation was first proposed in [31] to find estimates of mixing times in total variation distance. Further, one of the first Wasserstein bounds generated on a Markov chain model used CRN in its construction [20]. More recently, [5] showed how CRN simulation could be used for estimating an upper bound on the Wasserstein distance (their Proposition 3.1), and provided useful applications of the CRN method to high-dimensional and tall data (their Section 4). CRN simulation to estimate convergence rates has been used on a wide range of applications including Langevin algorithms [17], stochastic gradient descent [34], stochastic differential equations (SDE) [3], and Hamiltonian Monte Carlo [7, 22].
@@
The use of CRN to generate tight bounds on the Wasserstein and total variation distance is promising. In this paper we show that for the Gibbs sampler for a model related to James-Stein estimators and the variance component model, CRN bounds perform significantly better than drift and minorization (DnM) bounds (see figures 2 and 5 and commentary in Numerical Example 5.1 for more details). Under certain conditions, [3] and [21] show that CRN is the optimal coupling of the Wasserstein distance between simulated solutions of SDEs and two random variables in the $\mathcal{L}^{2}$ metric, respectively. [15] shows that CRN simulation on the independent Metropolis-Hastings algorithm is the maximal coupling of the total variation distance. [9] shows that the distance between two reflected Brownian motion processes simulated using CRN converge to 0 almost surely.
@@
In Theorem 3.2, using CRN simulation, we provide an estimated upper bound between the Wasserstein distance of a Markov chain and the corresponding stationary distribution when only the unnormalized density of the stationary distribution is known. Furthermore the $95\%$ confidence intervals of our estimates in the real data examples are fairly narrow. See Lemma 3.3 and its implementation in the figures 1, 4, 5 and 6 and tables 6 and 8 for more details.
@@
This paper is organized as follows. In Section 2, we present definitions and notation. In Section 3, we establish convergence bounds of a Markov chain to its corresponding stationary distribution using CRN simulation when the initial distribution is not in stationarity. In Section 4, we apply Theorem 3.2 to a Gibbs sampler for a model related to the James-Stein estimator and compare the CRN bound to a DnM and one-shot coupling bound in Real Data Example 4.1. In Section 5 we apply Theorem 3.2 to a Gibbs sampler for a variance component model and compare the CRN bound to a DnM bound in Numerical Example 5.1 and another DnM bound in Real Data Example 5.1. In Section 6 we apply Theorem 3.2 to a Gibbs sampler for a Bayesian linear regression model.
For each numerical example, we also provide the Gelman-Rubin diagnostic and traceplots for comparison. While the traceplots all started using the same initial distributions as our numerical examples, the Gelman-Rubin diagnostics are calculated from an initial distribution that is overdispersed with respect to the stationary distribution, so may not be perfectly comparable. Finally, while the Gelman-Rubin statistic and traceplots aid in determining convergence, they do not prove that convergence has occurred. Our method, on the other hand, provides an unbiased estimate of an upper bound on the Wasserstein distance between the distribution of our Markov chain and the stationary distribution.
@@ 2 Background and Notation
Let $(\mathcal{X},d)$ be a Polish metric space and $\mathcal{F}$ the associated Borel $\sigma$-algebra. The $L^{p}$-Wasserstein distance between distributions $\pi$ and $\nu$ on $(\mathcal{X},\mathcal{F})$ is defined as $W_{d,p}(\pi,\nu)=\inf_{X\sim\pi,Y\sim\nu}E[d(X,Y)^{p}]^{1/p}$. The Wasserstein space $P_{p}(\mathcal{X})$ is the set of Borel probability measures on the metric space $(\mathcal{X},d)$ such that for any $\mu\in P_{p}(\mathcal{X})$, $\int_{\mathcal{X}}d(x_{0},x)^{p}\mu(\mathrm{d}x)<\infty$ for some $x_{0}\in\mathcal{X}$. Total variation distance is defined as $\pi-\nu||_{TV}=\inf_{X\sim\pi,Y\sim\nu}P(X\neq Y)=\sup_{A\in\mathcal{F}}|\pi(A)-\nu(A)|$. The total variation distance can be written as a special case of the Wasserstein distance, $\pi-\nu||_{TV}=W_{d_{I},1}(\pi,\nu)$ where $d_{I}(X,Y)=I_{X\neq Y}$. If $\mathcal{X}\subset\mathbb{R}^{q}$, $q\in\mathbb{N}$, define the norm over a vector as $x||_{p}=(\sum_{i=1}^{q}|x_{i}|^{p})^{1/p}$ where $p\in[1,\infty)$. Define the p-norm over a function $\phi:\mathcal{X}\to\mathbb{R}$ as $\left\lVert\phi\right\rVert_{p}=\left(\int_{\mathcal{X}}\phi(x)^{p}\mathrm{d}x\right)^{\frac{1}{p}}$. If $X\perp\!\!\!\perp Y$, then $X$ and $Y$ are independent. If $\pi\ll\nu$, then $\pi$ is absolutely continuous with respect to $\nu$.
@@
Define a Markov chain $\{X_{n}\}_{n\geq 1}$ in $\mathcal{X}$ such that $X_{n}=k_{\theta_{n}}(X_{n-1})$, where $\{\theta_{n}\}_{n\geq 1}$ are i.i.d. random variables on some measurable space $\Theta$ and random measurable mappings $k_{\theta_{n}}:\mathcal{X}\mapsto\mathcal{X}$. The set of random functions $k_{\theta_{1}},k_{\theta_{2}},\ldots$ is called an iterated function system. Any time-homogeneous Markov chain can be represented as an iterated function system [53]. We can also write $X_{n}=k_{\theta_{n}}(k_{\theta_{n-1}}(k_{\theta_{n-2}}\ldots k_{\theta_{1}}(x_{0})))=k_{\Theta_{n}}(x_{0})$ where $\Theta_{n}=(\theta_{1},\ldots,\theta_{n})$ and $\theta_{i}\perp\!\!\!\perp\theta_{j}$, $i\neq j$. CRN simulation in this context occurs when we simulate two Markov chains $\{X_{n}\}_{n\geq 1}$ and $\{Y_{n}\}_{n\geq 1}$ with different initial values $x_{0}\neq y_{0}$, but with the same transition functions. That is, $\theta_{n}$ are the same in $X_{n}=k_{\theta_{n}}(X_{n-1})$ and $Y_{n}=k_{\theta_{n}}(Y_{n-1})$, $n\geq 1$. In different notation, the $\Theta_{n}$ is the same in $X_{n}=k_{\Theta_{n}}(x_{0})$ and $Y_{n}=k_{\Theta_{n}}(y_{0})$. See Algorithm 1 for more details. The stationary distribution of $X_{n}$ is defined as $\mathcal{L}(X_{\infty})=\pi$.
@@
$IG(\alpha,\beta)$ represents the inverse gamma distribution with density function $f_{IG}(x\mid\alpha,\beta)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}x^{-\alpha-1}e^{-\beta/x}$. $N(\mu,\sigma^{2})$ represents the normal distribution with density function $f_{N}(x\mid\mu,\sigma^{2})=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}$. For a distribution $\nu$, the density function is denoted as $f_{\nu}$.
3 Estimating Markov chain convergence rates using CRN simulation
We propose bounding the $L^{p}$-Wasserstein distance between the $n$th iteration of a Markov chain $\{X_{n}\}_{n\geq 1}$ and the corresponding stationary distribution $\pi$ through CRN simulation. Our main result is Theorem 3.2. Using CRN simulation as a convergence diagnostic tool was first discussed in [31] for bounding total variation distance.
We first provide a method for bounding the Wasserstein distance between $\mathcal{L}(X_{n})$ and the corresponding stationary distribution $\pi$, $W_{d,p}(\mathcal{L}(X_{n}),\pi)$ by proposing an auxiliary random variable $Y_{n}$. That is $W_{d,p}(\mathcal{L}(X_{n}),\pi)$ is bounded above by the product of the ‘distance’ between $\pi$ and $\mathcal{L}(Y_{n})$ (K) and $E[d(X_{n},Y_{n})]$.
@@ 3 Estimating Markov chain convergence rates using CRN simulation
Let $\{X_{n}\}_{n\geq 1}$ and $\{Y_{n}\}_{n\geq 1}$ be two copies of a Markov chain with stationary distribution $\pi$ and initial distributions $\mathcal{L}(Y_{0})=\nu$, $\mathcal{L}(X_{0})=\mu$ where $X_{0}\perp\!\!\!\perp Y_{0}$. Assume $\pi<<\nu$. Then for $r\geq p\geq 1$ and $s>1$, the following holds where $K=E\left[\left(\frac{f_{\pi}(Y_{0})}{f_{\nu}(Y_{0})}\right)^{\frac{s}{s-1}}\right]^{\frac{s-1}{s}}$
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq K^{1/r}E\left[d(X_{n},Y_{n})^{rs}\right]^{\frac{1}{rs}}$ (1)
See Section 8.2 for a proof.
Remark 1.
If we take $K=\lim_{s\downarrow 1}E\left[\left(\frac{f_{\pi}(Y_{0})}{f_{\nu}(Y_{0})}\right)^{\frac{s}{s-1}}\right]^{\frac{s-1}{s}}$, then from Theorem 3.1 we get that $K=ess\sup_{x\in\mathcal{X}}f_{\pi}(x)/f_{\nu}(x)$ and
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq K^{1/r}E\left[d(X_{n},Y_{n})^{r}\right]^{1/r}$
The case of $s\downarrow 1$ can be proven using the rejection sampler or the separation distance (see the proof of this special case in Section 8.3.1). The use of rejection sampling to generate an upper bound on the expected distance between a proposal density function $\nu$ and a stationary density function $\pi$ was motivated by [31], which used rejection sampling to generate similar upper bounds in total variation distance. This case is used in Lemmas 4.1 and 6.2
Remark 2.
Applying $d_{I}(X,Y)=I_{X\neq Y}$ to Remark 1, and setting $p,r=1$, the total variation distance between a Markov chain and the corresponding stationary distribution is bounded above as follows,
$\mathcal{L}(X_{n})-\pi||_{TV}\leq KP(X_{n}\neq Y_{n})$ (2)
And so, for stopping time $\tau=\min\{n:d(X_{n},Y_{n})=0\}$, $\mathcal{L}(X_{n})-\pi||_{TV}\leq KP(\tau>n)$.
In many cases, the normalizing constant for $f_{\pi}$ is unknown. That is, we only know of a function $g(x)$ such that $\frac{1}{L}g(x)=f_{\pi}(x)$ where $L$ is unknown. For cases like this, we note that for all $x\in\mathcal{X}$ and any $B\subset\mathcal{X}$,
$f_{\pi}(x)=\frac{g(x)}{\int_{\mathcal{X}}g(x)\mathrm{d}x}\leq\frac{g(x)}{\int_{B}g(x)\mathrm{d}x}$ (3)
This upper bound on $f_{\pi}$ is used to upper bound $K$ in all of the examples.
Next we define Algorithm 1, which generates an estimate of $E[d(X_{n},Y_{n})^{rs}]$ using CRN simulation.
Algorithm 1 An estimate using CRN of $E[d(X_{N},Y_{N})^{rs}]\approx\frac{1}{M}\sum_{i=1}^{M}d(x_{N,i},y_{N,i})^{rs}$
@@
Let $\{X_{n}\}_{n\geq 1}$ and $\{Y_{n}\}_{n\geq 1}$ be two copies of a Markov chain with stationary distribution $\pi$, initial distributions $\mathcal{L}(Y_{0})=\nu$, $\mathcal{L}(X_{0})=\mu$ ($X_{0}\perp\!\!\!\perp Y_{0}$) and simulated according to Algorithm 1. Assume $\pi<<\nu$. Then for $r\geq p\geq 1$ and $s\geq 1$, the following holds almost surely for $N\geq 1$, where $K=E\left[\left(\frac{f_{\pi}(Y_{0})}{f_{\nu}(Y_{0})}\right)^{\frac{s}{s-1}}\right]^{\frac{s-1}{s}}$ and $\mathcal{L}(X_{N}),\mathcal{L}(Y_{N})\in P_{rs}(\mathcal{X})$.
$W_{d,p}(\mathcal{L}(X_{N}),\pi)\leq K^{1/r}\left(\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}d(X_{N,i},Y_{N,i})^{rs}\right)^{\frac{1}{rs}}$ (4)